OpenStudy (anonymous):

Find the equivalent discount of 20 consecutive discounts of 10%

OpenStudy (anonymous):

If x is the total amount then \(\Hugex \times \left(\frac{(100-x)}{100}\right)^{20} \)

OpenStudy (anonymous):

\[ \Huge x \times \left(\frac{100-x}{100}\right)^{20} \]

OpenStudy (rulnick):

0.9^20 = 0.1216, so the equivalent discount is 1-0.1216 = 0.8784, or 87.84%.